OpenStudy (anonymous):

CALCULUS LIMIT: lim x->0.5- (2x-1)/|2x^3-x^2| I used direct substitution and gave me 0, but the answer says it's -4, why?

OpenStudy (anonymous):

show mw your work

OpenStudy (anonymous):

(2*0.5-1) / |2*0.5^3-0.5^2| = 0

hartnn (hartnn):

the denominator is also 0, so its 0/0 not only 0

OpenStudy (anonymous):

-(2x-1)/ x^2(2x-1)=- 1/x^2 so 1/(1/2^2)= -1/(1/4) so flip the fraction and - 4

OpenStudy (anonymous):

when doing calculus always reduce where you can

OpenStudy (anonymous):

do you see it the answer and why?

OpenStudy (anonymous):

yes thanks guy i got it now

OpenStudy (anonymous):

glad to help